OpenStudy (sammie):

a teacher asks the student to take any odd number, multiply by 4, add 16, and divide results by 2. subtract 7 from quotient and tell the answer. the teacher said she will be able to tell the student the original number. how?

OpenStudy (anonymous):

Let x = an odd number and y=answer. Then the equation you just described would look like this: y= ((x*4+16)/2)-7 If I told the teacher that my answer was 7, she could work backwards and find my original number: x=(2(y+7)-16)/4 x=(2(7+7)-16)/4 x=(2(14)-16)/4 x=(28-16)/4 x=12/4 x=3 No matter what your answer will be she can use the second equation I wrote to find your original odd number.